OpenStudy (anonymous):

How do you solve this problem? (3/4-2/3) + (1/2 + 1/3) = ? Common Denominator?

OpenStudy (anonymous):

(3/4 - 2/3) ... common denominator = 12 9/12 - 8/12 = 1/12 (1/2 + 1/3) ... common denominator = 6 3/6 + 2/6 = 5/6 1/12 + 5/6 = ..... common denominator = 12 1/12 + 10/12 = 11/12

OpenStudy (anonymous):

i got 12 on the first on n 6 on the second? How do u bring the 2 together?? Do you add them as it says or what?

OpenStudy (anonymous):

Where did u get 5/6? outta 1/2 n 1/3?

OpenStudy (anonymous):

Since 2 and 3 have no factors in common the common denominator = 2 * 3 = 6 You put 6 on the bottom of each fraction.. ???/6 + ???/6 then look at the first fraction which is 1/2 what do you need to multiply 2 by to get 6.. answer 3 so you multiply the top by 3 so 1 * 3 = 3 so that fraction is 3/6 Then look at the second fraction 1/3 what do you need to multiply 3 by to get 6.. answer 2 so you multiply the top by 2 so 1*2 = 2 so that fraction is 2/6 Then when you add fractions you don't touch the bottom, just add the tops so 3/6+ 2/6 = 5/6

OpenStudy (anonymous):

Then you need to add 1/12 + 5/6 ... the common denominator = 12 so put 12 on the bottom of each fraction. The first fraction already has a 12 on the bottom so it remains 1/12 to turn 5/6 into a 12 on the bottom you ask yourself, what do you multiply 6 by to get 12.. the answer is 2 so you multiply the top by 2 5*2 = 10 so the second fraction is 10/12 1/12 + 10/12 = 11/12 When you have common denominators, you don't touch the bottom just "add" the tops or subtract if that is the problem.